# World Models via Policy-Guided Trajectory Diffusion

### 13 Dec 2023 v1

@@ Abstract
  • World models are a powerful tool for developing intelligent agents. By predicting the outcome of a sequence of actions, world models enable policies to be optimised via on-policy reinforcement learning (RL) using synthetic data, i.e. in “in imagination”. Existing world models are autoregressive, and interleave predicting the next state with sampling the next action from the policy. Thus, the prediction error inevitably compounds as the trajectory length grows. In this work, we propose a novel world modelling approach that is not autoregressive and generates entire on-policy trajectories via a single pass through a diffusion model. Our approach, Policy-Guided Trajectory Diffusion (PolyGRAD), leverages a denoising model in addition to the gradient of the action distribution of the policy to diffuse a trajectory of initially random states and actions into an on-policy synthetic trajectory. We analyse the capabilities of our approach and demonstrate that it obtains competitive prediction errors to state-of-the-art autoregressive baselines. PolyGRAD also enables performant policies to be trained via on-policy RL in imagination. We believe that PolyGRAD introduces a promising paradigm for world modelling with many possible extensions to explore in future work.
@@ 1 Introduction
Model-based reinforcement learning (RL) trains a predictive model of the environment dynamics, conditioned upon actions. Such models are often referred to as world models. Once a world model has been learnt, imagined (i.e. synthetic) data generated from the world model can be utilising for planning (Schrittwieser et al., 2020) or on-policy RL (Hafner et al., 2021). By distilling the knowledge obtained about the environment into the world model, this approach facilitates improved sample efficiency relative to model-free approaches (Micheli et al., 2023) and zero-shot transfer to new tasks (Sekar et al., 2020).
@@
The core contributions of this work are: (a) proposing PolyGRAD, the first scalable approach to world modelling that enables on-policy trajectory generation without autoregressive sampling, and (b) analysing the connection between PolyGRAD, score-based generative models, and classifier-guided diffusion models. Our results demonstrate that PolyGRAD obtains prediction errors competitive with state-of-the-art autoregressive world models. We also show that PolyGRAD can be used to optimise performant policies via RL using only imagined data for standard continuous control benchmarks. Thus, PolyGRAD introduces a novel and promising paradigm for developing scalable world models without autoregressive sampling.
@@ 2 Related Work
There are several previous works that utilise diffusion models to generate synthetic data for RL. SynthER (Lu et al., 2023) trains a diffusion model to generate individual steps of off-policy synthetic data. This data is used to augment the replay buffer of an off-policy RL algorithm to improve sample efficiency. In a similar vein, MTDIFF-S (He et al., 2023) generates off-policy trajectories to augment offline RL datasets, but in the multi-task setting. Most related to our work are diffusion world models (Anonymous, 2023; Zhang et al., 2023), which train a diffusion model to generate a single step of transition data, conditioned on the current action and previous observations. Like other existing world models, this approach enables the generation of synthetic trajectories by sampling actions from the policy and autoregressively querying the model one step at a time. In contrast, our approach generates entire on-policy trajectories in a single pass of diffusion.
@@ 3 Preliminaries
Throughout this paper, we use subscript indices to refer to steps in a diffusion process and superscript indices to refer to time steps in a trajectory through the environment. Thus, $\mathbf{x}_{i}$ refers to a data sample at the $i^{\textnormal{th}}$ step in a diffusion process, while $s^{t}$ refers to a state at the $t^{\textnormal{th}}$ step in a trajectory. If a symbol has no subscript, it refers to a data sample that has had no noise added (i.e. $\mathbf{x}=\mathbf{x}_{0}$). Bold symbols refer to matrices.
@@ 4 World Models via Policy-Guided Trajectory Diffusion
In this work, we propose a new approach to world modelling: Policy-Guided tRAjectory Diffusion (PolyGRAD). A core novelty of PolyGRAD is that it is not an autoregressive world model that chains together a sequence of one-step predictions. Instead, PolyGRAD enables the generation of entire on-policy trajectories in a single pass through a diffusion model. The main challenge for creating a non-autoregressive world model is ensuring that the trajectories generated are sampled according to the current policy. To address this challenge, PolyGRAD utilises diffusion to gradually diffuse a trajectory of initially random states and actions into an on-policy synthetic trajectory.
@@ 4.1 Denoising Model Training (Algorithm 1)
Figure 2: Step-by-step llustration of PolyGRAD trajectory generation (Algorithm 2). Bottom left: Illustration of policy action distribution throughout state space. Top left: Trajectory is initialised with random states and actions (Line 2 and 3). The initial state (dark purple) is sampled from the dataset (Line 4). Solid black arrows: Action sequence is updated according to score of policy conditioned on current state sequence (Line 10). Hollow black arrows: State sequence is updated conditioned on current actions (Line 13). Bottom right: Final trajectory is returned.
@@ 4.2 Policy-Guided Trajectory Diffusion (Algorithm 2)
We now present PolyGRAD, our algorithm for generating imagined on-policy trajectories for world modelling (Algorithm 2). Each of the key steps of the algorithm is illustrated in Figure 2. PolyGRAD begins with a trajectory of random states, rewards, and actions, $\widehat{\bm{\tau}}_{N}=(\widehat{\bm{\tau}}^{sr}_{N},\widehat{\bm{\tau}}^{a}_{N})$, sampled from a standard normal distribution (Lines 2 and 3). Note that the subscript $N$ refers to the step in the diffusion process. We also sample an initial state $s_{0}$ uniformly from $\mathcal{D}$. The trajectory is then iteratively refined over many diffusion steps. To condition the trajectory on the initial state from the dataset $s_{0}$, we perform inpainting (Lugmayr et al., 2022) by replacing the initial state in the trajectory by $s_{0}$ at each step of the diffusion process (Line 6). Generating rollouts branched from states in the dataset is standard practice in model-based RL (Janner et al., 2019). At each diffusion step, the denoising model is used to predict the noise added to the state and reward sequence, conditioned on the current action sequence (Line 7). The predicted noise is denoted by $\widehat{\bm{\epsilon}}$.
In all steps except for the final diffusion step, the action sequence is updated in Lines 9 and 10. To perform the action update, we first use the predicted noise $\widehat{\bm{\epsilon}}$ to compute a prediction of the fully-denoised state sequence, $\widehat{\bm{\tau}}^{s}_{0}$. We compute this prediction because the policy is trained on data without noise added, so we do not wish to condition the policy on a noisy state sequence. In Line 10, we condition the policy on the predicted denoised state sequence. We update the action sequence in the direction of the score of the policy action distribution, $\nabla_{a}\pi_{\phi}(a|s)$, in Line 10. Note that this gradient differs from the standard gradient used in policy-gradient RL algorithms (which is $\nabla_{\phi}\pi_{\phi}(a|s)$), and can be easily computed for common policy parameterisations, such as Gaussian policies. The magnitude of the update to the actions is controlled by the action update scale, $\delta$, which as we shall see in Section 4.3, is tuned automatically online. In Line 13 the state sequence is updated using the noise prediction and a standard diffusion update (Equation 1). After completing all the diffusion steps, the final synthetic trajectory is returned in Line 14.
@@
6:   set initial state in $\widehat{\bm{\tau}}_{i}^{sr}$ to $s_{0}$triangleright Condition trajectory on initial state from dataset
7:   $\widehat{\bm{\epsilon}}\leftarrow\epsilon_{\theta}(\widehat{\bm{\tau}}^{sr}_{i}\ |\ i,\widehat{\bm{\tau}}^{a}_{i})$triangleright Predict state noise
@@
13:   $\widehat{\bm{\tau}}_{i-1}^{sr}\leftarrow\frac{1}{\sqrt{1-\beta_{i}}}(\widehat{\bm{\tau}}_{i}^{sr}-\frac{\beta_{i}}{\sqrt{1-\alpha_{i}}}\widehat{\epsilon}\ )+\sqrt{\beta}_{i}\mathbf{z}$triangleright Denoise states
@@
When $\beta\rightarrow 0$ and $t\rightarrow\infty$, stochastic Langevin dynamics converges to a sample from $p(\mathbf{x})$ (Welling & Teh, 2011; Song & Ermon, 2019). Score-based generative models attempt to learn an approximation to the score function, $s_{\theta}(\mathbf{x})\approx\nabla_{\mathbf{x}}\log p(\mathbf{x})$, and then perform sampling according to Equation 5 using $s_{\theta}$. As noted in Section 3, diffusion models are in instance of score-based generative models, where the score of the noisy data distribution is learned across a range of noise levels.
@@ 5 Experiments
In our experiments, we seek to answer the following questions: (a) Does PolyGRAD produce the correct action distribution? (b) How accurate are the trajectories produced by PolyGRAD? (c) Can the synthetic data produced by PolyGRAD be used to train performant policies? (d) Which implementation details influence the performance of PolyGRAD? To answer these questions, we run experiments using the MuJoCo environments in OpenAI Gym (Brockman et al., 2016). The code for our experiments is available at github.com/marc-rigter/polygrad-world-models.
@@ 5.1 Does PolyGRAD produce the correct action distribution?
We wish to evaluate whether PolyGRAD generates synthetic trajectories with the correct on-policy action distribution. To evaluate the action distribution produced, we train an RL policy using synthetic data generated by PolyGRAD (Algorithm 3). We linearly decay the standard deviation of the Gaussian policy, $\sigma_{\phi}$, from 1 to 0.02 over the course of training. In Figure 3, we plot the distribution of the difference between the action $a$ and the mean of the action distribution, $\mu_{\phi}(s)$, over all state-action pairs $(s,a)$ in batches of data produced by PolyGRAD with $h=10$.
We observe that for $\sigma_{\phi}\geq 0.1$, the action distribution closely matches the target Gaussian distribution. This demonstrates that PolyGRAD produces the correct action distribution on aggregate provided that the policy entropy is not too low. However, we observe that as the policy standard deviation decreases below 0.1, the action distribution begins to deviate from the target distribution. Thus, if the policy has very low entropy, it is difficult for PolyGRAD to guide the initially randomly sampled actions to the target distribution. In the plot with $\sigma_{\phi}=0.02$, the distribution has additional modes at ±$\pm 3$ standard deviations. This is an artifact of the action clipping described in the implementation details. Equivalent plots for the other MuJoCo environments are provided in Appendix B.2 of the supplementary material. Appendix B.3 shows that the action distributions produced are similar for longer trajectories with $h=50$.
@@
Figure 3: Plots of action distributions produced by PolyGRAD. Blue line illustrates the distribution of $a-\mu_{\phi}(s)$ for a batch of synthetic data. Each subplot is for a policy with a different entropy level that is constant throughout the state space. Dashed black line indicates the action distribution output by the policy. Data is generated by running Algorithm 3 in Walker2d with $h=10$.
5.2 How accurate are the trajectories produced by PolyGRAD?
To assess the accuracy of the trajectories produced by PolyGRAD, we train several different world models using the same dataset of 1M transitions in each environment. We then generate synthetic rollouts using the same performant policy in each world model. We replay the same actions in the MuJoCo simulator, and evaluate the errors in the trajectory predictions. We compare the accuracy of the trajectories against the following baselines:
@@ 5.2 How accurate are the trajectories produced by PolyGRAD?
Figure 5 shows the error for each world modelling method for varying trajectory lengths. We train PolyGRAD for trajectories of $h=10,50$, and $200$. For $h=10$, PolyGRAD obtains the equal-best trajectory errors for Hopper and Walker2d along with Autoregressive Diffusion. For HalfCheetah, PolyGRAD with $h=10$ obtains the second-best errors to Autoregressive Diffusion. When PolyGRAD is trained on longer trajectories ($h=50$ and $h=200$) we observe that larger errors are obtained. This indicates that it is more challenging for the denoising model to accurately predict the denoised states for longer trajectories. The errors for the Probabilistic MLP Ensemble increase quickly as the prediction horizon increases. This may be because this method outputs a Gaussian distribution over the next state, and repeatedly sampling from this Gaussian can quickly lead to out-of-distribution states.
@@
The Transformer baseline obtains comparable performance to PolyGRAD with $h=200$, but PolyGRAD outperforms the transformer when trained and evaluated on shorter trajectories. Note that PolyGRAD uses the same transformer network architecture for the denoising model as the Transformer baseline. Therefore, this result demonstrates that there is an advantage to iteratively refining the trajectory via diffusion rather than autoregressively predicting each successor state with a transformer.

### 17 Dec 2023 v2

@@ Abstract
  • World models are a powerful tool for developing intelligent agents. By predicting the outcome of a sequence of actions, world models enable policies to be optimised via on-policy reinforcement learning (RL) using synthetic data, i.e. in “in imagination”. Existing world models are autoregressive in that they interleave predicting the next state with sampling the next action from the policy. Prediction error inevitably compounds as the trajectory length grows. In this work, we propose a novel world modelling approach that is not autoregressive and generates entire on-policy trajectories in a single pass through a diffusion model. Our approach, Policy-Guided Trajectory Diffusion (PolyGRAD), leverages a denoising model in addition to the gradient of the action distribution of the policy to diffuse a trajectory of initially random states and actions into an on-policy synthetic trajectory. We analyse the connections between PolyGRAD, score-based generative models, and classifier-guided diffusion models. Our results demonstrate that PolyGRAD outperforms state-of-the-art baselines in terms of trajectory prediction error for moderate-length trajectories, with the exception of autoregressive diffusion. At short horizons, PolyGRAD obtains comparable errors to autoregressive diffusion, but with significantly lower computational requirements. Our experiments also demonstrate that PolyGRAD enables performant policies to be trained via on-policy RL in imagination for MuJoCo continuous control domains. Thus, PolyGRAD introduces a new paradigm for scalable and non-autoregressive on-policy world modelling.
@@ 1 Introduction
Model-based reinforcement learning (RL) trains a predictive model of the environment dynamics, conditioned upon actions. Such models are often referred to as world models. Once a world model has been learnt, imagined (i.e. synthetic) data generated from the world model can be utilised for planning (Schrittwieser et al., 2020) or on-policy RL (Hafner et al., 2021). By distilling the knowledge obtained about the environment into the world model, this approach facilitates improved sample efficiency relative to model-free approaches (Micheli et al., 2023) and zero-shot transfer to new tasks (Sekar et al., 2020).
@@
The core contributions of this work are: (a) proposing PolyGRAD, the first scalable approach to world modelling that enables on-policy trajectory generation without autoregressive sampling, and (b) analysing the connection between PolyGRAD, score-based generative models, and classifier-guided diffusion models. Our results demonstrate that PolyGRAD outperforms state-of-the-art baselines in terms of trajectory prediction error for moderate-length trajectories, with the exception of autoregressive diffusion. At short horizons, PolyGRAD obtains comparable errors to autoregressive diffusion, but with significantly lower computational requirements. We also show that PolyGRAD can be used to optimise performant policies via on-policy RL using only imagined data for MuJoCo continuous control domains. Thus, PolyGRAD introduces a new paradigm for developing scalable and accurate world models without autoregressive sampling.
@@ 2 Related Work
There are several previous works that use diffusion models to generate synthetic data for RL (Zhu et al., 2023). SynthER (Lu et al., 2023) trains a diffusion model to generate individual steps of off-policy synthetic data. This data is used to augment the replay buffer of an off-policy RL algorithm to improve sample efficiency. Similarly, MTDIFF-S (He et al., 2023) generates off-policy trajectories to augment offline RL datasets, but in the multi-task setting. Concurrent works on world modelling with diffusion (Anonymous, 2023; Zhang et al., 2023) train a diffusion model to generate a single step of transition data, conditioned on the current action and previous observations. Like other existing world models, this approach enables the generation of synthetic trajectories by sampling actions from the policy and autoregressively querying the model one step at a time. In contrast, our approach generates entire on-policy trajectories in a single pass of diffusion. Also concurrently, and most related to our work, Jackson et al. (2023) guide diffusion with a policy to increase the likelihood of generated trajectories under the policy, in a similar manner to PolyGRAD. Jackson et al. (2023) use this approach to generate synthetic datasets for offline, off-policy RL. In contrast, PolyGRAD generates on-policy trajectories for online, on-policy RL in imagination. Unlike Jackson et al. (2023), we also analyse the connection between PolyGRAD, classifier-guidance, and score-based generative models. Furthermore, we show that the scale of the action guidance can be tuned automatically online to approximately generate the correct on-policy action distribution.
@@ 3 Preliminaries
Throughout this paper, we use subscript indices to refer to steps in a diffusion process and superscript indices to refer to time steps in a trajectory through the environment. Thus, $\mathbf{x}_{i}$ refers to a data sample at the $i^{\textnormal{th}}$ step in a diffusion process, while $a^{t}$ refers to an action at the $t^{\textnormal{th}}$ step in a trajectory. If a symbol has no subscript, it refers to a data sample that has had no noise added (i.e. $\mathbf{x}=\mathbf{x}_{0}$). Bold symbols refer to matrices.
@@ 4 World Models via Policy-Guided Trajectory Diffusion
In this work, we propose a new approach to world modelling: Policy-Guided tRAjectory Diffusion (PolyGRAD). A core novelty of PolyGRAD is that it enables the generation of entire on-policy trajectories in a single pass of diffusion, rather than autoregressively chaining a sequence of one-step predictions. The main challenge for creating a non-autoregressive world model is ensuring that the trajectories generated are sampled according to the current policy. To address this challenge, PolyGRAD utilises diffusion to gradually diffuse a trajectory of initially random states and actions into an on-policy synthetic trajectory.
@@ 4.2 Policy-Guided Trajectory Diffusion (Algorithm 2)
We now present PolyGRAD, our algorithm for generating imagined on-policy trajectories for world modelling (Algorithm 2). Each of the key steps of the algorithm is illustrated in Figure 2. PolyGRAD begins with a trajectory of random states, rewards, and actions, $\widehat{\bm{\tau}}_{N}=(\widehat{\bm{\tau}}^{sr}_{N},\widehat{\bm{\tau}}^{a}_{N})$, sampled from a standard normal distribution (Lines 2 and 3). Note that the subscript $N$ refers to the step in the diffusion process. We also sample an initial state $s^{0}$ uniformly from $\mathcal{D}$. The trajectory is then iteratively refined over many diffusion steps. To condition the trajectory on the initial state from the dataset $s^{0}$, we perform inpainting (Lugmayr et al., 2022) by replacing the initial state in the trajectory by $s^{0}$ at each step of the diffusion process (Line 6). Generating rollouts branched from states in the dataset is standard practice in model-based RL (Janner et al., 2019). At each diffusion step, the denoising model is used to predict the noise added to the state and reward sequence, conditioned on the current action sequence (Line 7). The predicted noise is denoted by $\widehat{\bm{\epsilon}}$.
In all steps except for the final diffusion step, the action sequence is updated in Lines 9 and 10. To perform the action update, we first use the predicted noise $\widehat{\bm{\epsilon}}$ to compute a prediction of the fully-denoised state sequence, $\widehat{\bm{\tau}}^{s}_{0}$. We compute this prediction because the policy is trained on data without noise added, so we do not wish to condition the policy on a noisy state sequence. In Line 10, we condition the policy on the predicted denoised state sequence. We update the action sequence in the direction of the score of the policy action distribution, $\nabla_{a}\pi_{\phi}(a|s)$, in Line 10. Note that this gradient differs from the standard gradient used in policy-gradient RL algorithms (which is $\nabla_{\phi}\pi_{\phi}(a|s)$), and can be easily computed for common policy parameterisations, such as Gaussian policies. The magnitude of the update to the actions is controlled by the action update scale, $\delta$, which as we shall see in Section 4.3, is tuned automatically online.
Figure 2: Step-by-step illustration of PolyGRAD trajectory generation (Algorithm 2). Bottom left: Illustration of policy action distribution throughout state space. Top left: Trajectory is initialised with random states and actions (Line 2 and 3). The initial state (dark purple) is sampled from the dataset (Line 4). Solid black arrows: Action sequence is updated according to score of policy conditioned on current state sequence (Line 10). Hollow black arrows: State sequence is updated conditioned on current actions (Line 13). Bottom right: Final trajectory is returned.
In Line 13 the state and reward sequence is updated using the noise prediction and a standard diffusion update (Equation 1). After completing all the diffusion steps, the final synthetic trajectory is returned in Line 14.
@@
6:   set initial state in $\widehat{\bm{\tau}}_{i}^{sr}$ to $s^{0}$triangleright Condition trajectory on initial state from dataset
7:   $\widehat{\bm{\epsilon}}\leftarrow\epsilon_{\theta}(\widehat{\bm{\tau}}^{sr}_{i}\ |\ i,\widehat{\bm{\tau}}^{a}_{i})$triangleright Predict noise
@@
13:   $\widehat{\bm{\tau}}_{i-1}^{sr}\leftarrow\frac{1}{\sqrt{1-\beta_{i}}}(\widehat{\bm{\tau}}_{i}^{sr}-\frac{\beta_{i}}{\sqrt{1-\alpha_{i}}}\widehat{\epsilon}\ )+\sqrt{\beta}_{i}\mathbf{z}$triangleright Update states and rewards using diffusion (Equation 1)
@@
When $\beta\rightarrow 0$ and $t\rightarrow\infty$, stochastic Langevin dynamics converges to a sample from $p(\mathbf{x})$ (Welling & Teh, 2011; Song & Ermon, 2019). Score-based generative models attempt to learn an approximation to the score function, $s_{\theta}(\mathbf{x})\approx\nabla_{\mathbf{x}}\log p(\mathbf{x})$, and then perform sampling according to Equation 5 using $s_{\theta}$. As noted in Section 3, diffusion models are in instance of score-based generative models, where the score of the noisy data distribution is learnt across a range of noise levels.
@@ 5 Experiments
In our experiments, we seek to answer the following questions: (a) Does PolyGRAD produce the correct action distribution? (b) How accurate are the trajectories produced by PolyGRAD compared to autoregressive world models? (c) Can the synthetic data produced by PolyGRAD be used to train performant policies? (d) Which implementation details influence the performance of PolyGRAD? To answer these questions, we run experiments using the MuJoCo environments in OpenAI Gym (Brockman et al., 2016). The code for our experiments is available at github.com/marc-rigter/polygrad-world-models.
@@ 5.1 Does PolyGRAD produce the correct action distribution?
We wish to evaluate whether PolyGRAD generates synthetic trajectories with the correct on-policy action distribution. To evaluate the action distribution produced, we train an RL policy using synthetic data generated by PolyGRAD (Algorithm 3). We linearly decay the standard deviation of the Gaussian policy, $\sigma_{\phi}$, from 1 to 0.02 over the course of training. In Figure 3, we plot the distribution of the difference between the action $a$ and the mean of the action distribution, $\mu_{\phi}(s)$, over all state-action pairs $(s,a)$ in batches of data produced by PolyGRAD with $h=50$ in Walker2d.
We observe that for $\sigma_{\phi}\geq 0.1$, the action distribution closely matches the target Gaussian distribution. This demonstrates that PolyGRAD produces the correct action distribution on aggregate provided that the policy entropy is not too low. However, we observe that as the policy standard deviation decreases below 0.1, the action distribution begins to deviate from the target distribution. Thus, if the policy has very low entropy, it is difficult for PolyGRAD to guide the initially randomly sampled actions to the target distribution. In the plot with $\sigma_{\phi}=0.02$, the distribution has additional modes at ±$\pm 3$ standard deviations. This is an artifact of the action clipping described in the implementation details. Plots for the other MuJoCo environments are provided in Appendix B.2 of the supplementary material.
@@
Figure 3: Plots of action distributions produced by PolyGRAD. Blue line illustrates the distribution of $a-\mu_{\phi}(s)$ for a batch of synthetic data. Each subplot is for a policy with a different entropy level that is constant throughout the state space. Dashed black line indicates the action distribution output by the policy. Data is generated by running Algorithm 3 in Walker2d with $h=50$.
5.2 How accurate are the trajectories produced by PolyGRAD compared to autoregressive world models?
To assess the accuracy of the trajectories produced by PolyGRAD, we train several different world models using the same dataset of 1M transitions in each environment. We then generate synthetic rollouts using the same performant policy in each world model. We replay the same actions in the MuJoCo simulator, and evaluate the errors in the trajectory predictions. We compare the accuracy of the trajectories against the following baselines which all generate trajectories autoregressively:
@@ 5.2 How accurate are the trajectories produced by PolyGRAD compared to autoregressive world models?
Figure 5 shows the error for each world modelling method for varying trajectory lengths. We train PolyGRAD for trajectories of $h=10,50$, and $200$. For $h=10$, PolyGRAD obtains the equal-best trajectory errors for Hopper and Walker2d along with Autoregressive Diffusion. For HalfCheetah, PolyGRAD with $h=10$ obtains the second-best errors to Autoregressive Diffusion. When PolyGRAD is trained on longer trajectories ($h=50$ and $h=200$) we observe that larger errors are obtained. This indicates that it is more challenging for the denoising model to accurately predict the denoised states for longer trajectories.
@@
The errors for the Probabilistic MLP Ensemble increase quickly as the prediction horizon increases. This may be because this method outputs a Gaussian distribution over the next state, and repeatedly sampling from this Gaussian can quickly lead to out-of-distribution states. The Transformer baseline obtains comparable performance to PolyGRAD with $h=200$, but PolyGRAD outperforms the transformer when trained and evaluated on shorter trajectories. Note that PolyGRAD uses the same transformer network architecture for the denoising model as the Transformer baseline. Therefore, this result demonstrates that there is an advantage to iteratively refining the trajectory via diffusion rather than autoregressively predicting each successor state with a transformer.